Rehabilitation Telecommunication System

Northeast Parallel Architectures Center Syracuse University

Statement of Work

We will develop multimedia (video/image/audio/text) on-demand databases for blind and visually impaired users. We will utilize the high-speed computing and network resources at NPAC to develop database servers that can be accessed remotely via NYNet. The servers will utilize the nCUBE parallel computer and high-end workstations. The multimedia databases will include digitized videos of rehabilitation instruction sessions; a catalog of available low-vision aids with accompanying descriptions (images, text and audio); information about available programs and benefits for the blind and visually impaired (text and audio); and a browsable library of audio books.

The above services can be delivered using Mosaic, a network multimedia browser. However Mosaic has an advanced graphical user interface that is very visually orientated, so is not suitable for the blind or visually impaired. Other network information tools such as gopher are easily usable by the blind through a text-to-speech translating device, since the information is completely text based. We will develop a Mosaic interface that is more suitable for blind or low-vision users.